OpenStudy (grace111):

Which is more ? 50% of 36 or 25 % of 60 ?

OpenStudy (vshiroky):

Multiply the 2 36*.5 50*.25

OpenStudy (anonymous):

first 50 % of 36 = 36 * 50/100 =36 * 0.5 = 18 and similary for 25 % of 60

OpenStudy (anonymous):

not divide multiply

OpenStudy (timmyk14):

You know I just remembered that

OpenStudy (timmyk14):

Okay, let me do this again lol

OpenStudy (anonymous):

for 25 % 60 we mean that 60 x 25/100 60 x 0.25 15 so which would greater 15 or 18

OpenStudy (grace111):

are you asking which is greater 15 or 18?

OpenStudy (anonymous):

ya u get the answer by the steps i hope so ........

OpenStudy (grace111):

so its 18

OpenStudy (anonymous):

yup or we call as 50% of 36

OpenStudy (grace111):

oh thanks ;)

OpenStudy (anonymous):

it's ok ....(^_^)